OpenStudy (anonymous):

Find the missing side length of each right triangle. Round to the nearest tenth. 1.a= 12,b= 9,c= ? 2.a= 14,b= ?,c= 35 3.a= ?,b= 4,c= 41

OpenStudy (anonymous):

the formula used here is... pythagoras theorem.: a^2+b^2=c^2 1.sqrt(63)=7.937=7.94 u get it? as in , how to round off?